# Interpretation of 95 GeV Excess within the Georgi-Machacek Model in Light of Positive Definiteness Constraints

### v1

@@ 1 Introduction
$\mu_{\gamma\gamma}^{ATLAS+CMS}={0.24}_{-0.8}^{+0.09}.$ (1)
@@
Given such interesting coincidental signals, it is meaningful to explore whether the final result corresponds to a new light scalar particle from the perspective of new physics. Indeed, many papers have explained these signals within different new physics models combine:95 ; 95-1 ; 95-2 ; 95-3 ; 95-4 ; 95-5 ; 95-6 ; 95-7 ; 95-8 ; 95-9 ; 95-10 ; 95-11 ; 95-12 ; 95-13 ; 95-14 ; 95-15 ; 95-16 ; 95-17 ; 95-18 ; 95-19 ; 95-20 ; 95-21 ; 95-22 ; 95-23 ; 95-24 ; 95-25 ; 95-26 ; 95-27 ; 95-28 ; 95-29 ; 95-30 ; 95-31 ; 95-32 ; 95-33 ; 95-34 ; 95-35 ; 95-36 ; 95-37 ; 95-38 ; 95-39 ; 95-40 ; 95-GM-1 ; 95-GM-2 .
@@
Among these models, the GM model can provide a light custodial singlet scalar as a perfect candidate to interpret di-photon and $b\bar{b}$ excess simultaneously 95-GM-1 . There could even be a twin-peak resonance around 95 GeV 95-GM-2 , and demonstrates the nature of di-photon, $b\bar{b}$, and di-tau excesses with two scalars in the GM model (a light CP-even scalar and a same-mass CP-odd scalar).
@@
The GM model  95-GM-1 ; 95-GM-2 ; GM ; GM2 ; GM3 ; GM4 ; GM5 ; GM6 ; GM7 ; GM8 ; GM9 ; GM10 ; GM11 ; GM12 ; GM13 ; GM14 ; GM15 ; GM16 ; GM17 ; GM18 ; GM19 ; GM20 ; GM21 ; GM22 ; GM23 ; GM24 ; GM25 ; GM26 ; GM27 ; GM28 ; GM29 ; GM30 ; GM31 ; GM32 ; GM33 ; GM34 ; GM35 ; GM36 ; GM37 ; GM38 ; GM39 ; GM40 ; GM41 ; GM63 ; GM64 ; GM66 ; GM44 ; GM45 ; GM46 ; GM47 ; GM48 ; GM49 ; GM50 ; GM51 ; GM52 ; GM53 ; GM54 ; GM55 ; GM56 ; GM57 ; GM58 ; GM59 ; GM60 ; GM61 ; GM62 ; GM67 ; GM65 ; GM-pd extends the SM by incorporating a complex $SU(2)_{L}$ triplet scalar $\chi$ with unit hypercharge and a real $SU(2)_{L}$ triplet scalar $\xi$ with zero hypercharge. This intricate construction enables the nontrivial preservation of custodial symmetry at tree level after EWSB, dependent on the precise alignment of the vacuum expectation values (VEVs) between these triplet scalars, which ensures that the triplet state expectation value can reach several tens of GeV, significantly higher than in new physics models that extend the SM with a triplet. In addition to the unique additional scalar sector, the GM model predicts interesting phenomenology GM24 ; GM25 ; GM26 ; GM27 ; GM28 ; GM29 ; GM30 ; GM31 ; GM32 ; GM33 ; GM34 ; GM35 ; GM36 ; GM37 ; GM38 ; GM39 ; GM40 ; GM41 ; GM63 ; GM64 ; GM66 , such as enhanced couplings of the SM-like Higgs boson to W and Z gauge bosons, as well as the existence of nine more physical scalars, including a doubly charged scalar particle.
@@
In the GM model related studies, theoretical constraints from perturbative unitarity, the EW vacuum stability and conditions to avoid alternative minima vacuum GM57 ; GM58 ; GM59 ; GM60 ; GM61 ; GM62 ; GM67 etc should be satisfied other than those exclusion bounds from new particle searches GM56 by colliders. In our previous work GM-pd , we have rediscussed the vacuum stability bound for the GM model, concentrating specially on the constraints bounds from non-existence of a deeper vacuum at large field values. The numerical results indicate that quite a lot parameter space satisfied theoretical and experimental constraints have been excluded by the new vacuum stability bounds. It is interesting to explore whether there is still parameter space within the GM model that can accommodate the 95 GeV excess, considering all the theoretical and experimental constraints.
@@ 2 A brief review of the Georgi-Machacek model
The GM model’s Higgs sector encompasses the well-known SM-like $SU(2)_{L}$ complex doublet Higgs field, labeled as $({\phi}^{+},~{}{\phi}^{0})$, which carries a hypercharge of $1/2$. Furthermore, it incorporates two additional $SU(2)_{L}$ triplet Higgs fields: one complex triplet, identified as $({\chi}^{++},~{}{\chi}^{+},~{}{\chi}^{0})$) with a unit hypercharge, and a real triplet possessing a zero hypercharge, denoted as $({\xi}^{+},~{}{\xi}^{0},~{}-{\xi}^{+*})$. This extension introduces novel features and complexities in the Higgs sector compared to the Standard Model. The general potential [^2] The last two terms in the scalar potential can be absent if introducing a $Z_{2}$ symmetry. of GM model being invariant under the global $SU(2)_{L}\times SU(2)_{R}$ symmetry and SM gauge symmetry can be given as
@@
$m_{h,H}^{2}=\frac{\mathcal{M}_{11}^{2}+\mathcal{M}_{22}^{2}\mp\sqrt{{\left(\mathcal{M}_{11}^{2}-\mathcal{M}_{22}^{2}\right)}^{2}+4{\left(\mathcal{M}_{12}^{2}\right)}}}{2}.$ (48)
@@ 3.1 Brief review of the 95 GeV excess
$\mu_{\gamma\gamma}^{ATLAS+CMS}\equiv\frac{\sigma^{exp}(gg\rightarrow\phi\rightarrow\gamma\gamma)}{\sigma^{SM}(gg\rightarrow h\rightarrow\gamma\gamma)}={0.24}_{-0.8}^{+0.09}.$ (51)
@@ 3.2 The 95 GeV excess at the GM model
$\mu(\Phi_{95\mathrm{GeV}})\equiv\frac{\sigma({\Phi_{95\mathrm{GeV}}})\times BR(\Phi_{95\mathrm{GeV}}\rightarrow\gamma\gamma/b\bar{b})}{\sigma_{SM}({\Phi_{95\mathrm{GeV}}})\times BR_{SM}(\Phi_{95\mathrm{GeV}}\rightarrow\gamma\gamma/b\bar{b})}.$ (53)
@@
$\kappa_{Hff}=g_{H_{GM}ff}^{2}/g_{h_{SM}ff}^{2},~{}~{}~{}\kappa_{HVV}=g_{H_{GM}VV}^{2}/g_{h_{SM}VV}^{2}.$ (55)
In the investigation of resonances around 95 GeV, the choice for $ff$ should be $t\bar{t}$, and for $VV$, it should be $ZZ$. These selections correspond to the production modes that are relevant to both LEP and LHC experiments. For the GM model, the effective Higgs couplings to $t\bar{t}$ and $ZZ$ in the lowest order can be shown as
$\sqrt{\kappa_{Ht\bar{t}}}=\frac{\sin{\alpha}}{\cos{\theta}},~{}~{}~{}\sqrt{\kappa_{ht\bar{t}}}=\frac{\cos{\alpha}}{\cos{\theta}},$
$\sqrt{\kappa_{HZZ}}=\cos{\theta}\sin{\alpha}+\sqrt{\frac{8}{3}}\sin{\theta}\cos{\alpha},~{}~{}~{}\sqrt{\kappa_{hZZ}}=\cos{\theta}\cos{\alpha}-\sqrt{\frac{8}{3}}\sin{\theta}\sin{\alpha},$ (56)
@@ 4.1 The Usual Constraints
  • Relevant constraints about Higgs: The mass of the light singlet scalar $H$ is anticipated to be around 95.4 GeV to explain the observed di-photon and $b\bar{b}$ excesses, with a theoretical and experimental uncertainty of ±$\textpm 1$ GeV, setting the mass range from 94.4 GeV to 96.4 GeV. The properties of the heavier singlet scalar $h$, corresponding to the SM-like Higgs boson found at the LHC, should match ATLAS and CMS measurements at the $95\%$ confidence level, requiring p-value greater than 0.05 and $\Delta_{\chi^{2}}=\chi^{2}_{total}-\chi^{2}_{SM}$ smaller than 6.18, as verified by the package of HiggsSignal-2.6.2 HiggsSignals1 ; HiggsSignals2 ; HiggsSignals3 ; HiggsTools . Moreover, the signal rates for extra Higgs bosons must be consistent with cross-section limits from collider data (including LHC, LEP, etc), which can be checked by the package of HiggsBounds-5.10 HiggsBounds1 ; HiggsBounds2 ; HiggsBounds3 ; HiggsBounds4 .
@@
    As noted in eq(55), the couplings of SM-like Higgs $h$ to the SM fermions and weak gauge bosons $V=W,Z$ are modified in GM-type models, and the deviations from the SM predictions can be parameterized by the so-called “effective Higgs coupling”$\kappa_{ff}$ and $\kappa_{VV}$. The most strict constraints are given by ATLAS and CMS collaborations ATLAS:kappaWZ ; CMS:kappaWZ ; PDG:2022
@@
    $\kappa_{tt}=1.01\pm 0.11,~{}\kappa_{bb}=0.99\pm 0.16,~{}\kappa_{\tau\tau}=0.92\pm 0.08,$ (74)
@@
    $0.83\leq\kappa_{tt}\leq 1.10,~{}0.78\leq\kappa_{bb}\leq 1.15,~{}0.86\leq\kappa_{\tau\tau}\leq 1.00,$ (76)
@@
    It appears that the effective Higgs coupling $\kappa_{ff}$ tends to be less than 1, whereas the $\kappa_{VV}$ tend to be slightly greater than 1. This characteristic is a notable feature of the GM model and adds another intriguing layer to our understanding of it.
@@
    $1.7\times 10^{-9}<BR(B_{s}\rightarrow\mu^{+}\mu^{-})<4.5\times 10^{-9}~{},$ (78)
    $0.85\times 10^{-4}<BR(B^{+}\rightarrow\tau^{+}\nu)<2.89\times 10^{-4}~{},$ (79)
    $2.99\times 10^{-4}<BR(B_{S}\rightarrow X_{s}\gamma)<3.87\times 10^{-4}~{}.$ (80)
@@ 4.2 Positive Definiteness Constraints
In light of these considerations, we have introduced the so-called “Positive Definiteness Constraints” within the GM model to ensure the stability of the custodial vacuum. Here we briefly review the Positive Definiteness Constraints on the GM model.
@@
    $+\mu_{1}\phi^{\dagger}\xi\phi+\mu_{2}[\phi^{T}(i\tau_{2})\chi^{\dagger}\phi+\text{h.c.}]+\mu_{3}\text{tr}(\chi^{\dagger}\chi\xi)$
@@
    $m_{\phi}^{2}$ $=$ $2m_{\Phi}^{2},~{}m_{\chi}^{2}=2m_{\Delta}^{2},~{}m_{\xi}^{2}=m_{\Delta}^{2},~{}\mu_{1}=-\frac{M_{1}}{\sqrt{2}},~{}$
    $\mu_{2}$ $=$ $-\frac{M_{1}}{2},~{}\mu_{3}=6\sqrt{2}M_{2},\lambda=4\lambda_{1},~{}\rho_{1}=4\lambda_{2}+6\lambda_{3},~{}$
@@
    A comprehensive analysis of the effects of custodial symmetry breaking in the GM model at higher energies can be found in Ref. GM44 , assuming that no custodial symmetry-violating terms are present at the EW scale. The study demonstrates that, within certain parameter regions, the influence of custodial symmetry breaking can be adequately controlled up to the grand unified theory (GUT) scale. Alternatively, if custodial symmetry is preserved at high-energy scales within a UV completion of the GM model, the measured value of the electroweak $\rho$ parameter, along with the condition of perturbative unitarity, places stringent limits on the scale of this ultraviolet completion across almost the entire parameter space GM . Despite their different approaches, the two methods used to explore the custodial symmetry breaking effects induced by RGE between various energy scales yield equivalent consequences in the frame of the GM model.
    To conduct the RGE analysis, it is necessary to derive the $\beta$ functions for the couplings of the GM model, such as the scalar quartic couplings $\lambda,~{}\rho_{1,2,3,4,5},~{}\sigma_{1,2,3,4}$, the gauge couplings $g_{i}(i=1,2,3)$, and the Yukawa couplings $Y_{t,b,\tau}$, etc. Two-loop beta functions for a generic quantum field theory are documented in references Machacek:1983tz ; Machacek:1983fi ; Machacek:1984zw . For the GM model specifically, one-loop Renormalization Group Equations (RGEs) have been provided in references GM44 ; GM45 , with two-loop RGEs available in reference GM65 . In practical numerical analyses, the Mathematica package SARAH SARAH1 ; SARAH2 ; SARAH3 ; SARAH4 ; SARAH5 not only facilitates the generation of source codes for the spectrum-generator package SPheno SPheno1 ; SPheno2 but also can be utilized for RGE numerical studies. SARAH has included the GM model within its non-SUSY model database, which can simplify our study.
@@
    The invariance of the $SU(2)_{L}$ gauge symmetry allows us to eliminate certain redundant VEVs for one of the weak isospin components. Consequently, we can assume that $\phi^{+}=0$ at the minimum of the potential. Furthermore, the invariance of the $SU(2)_{V}$ custodial symmetry can be used to render three of the VEVs within the triplets redundant. If we choose to eliminate the VEVs for $\chi^{++}$ and assume CP conservation (where the imaginary parts of the VEVs vanish), we are left with the following potential form of the VEVs
    $\langle\phi\rangle=\left(\begin{array}[]{c}0\\ \frac{v_{\phi}}{\sqrt{2}}\end{array}\right),~{}\langle\chi\rangle=\left(\begin{array}[]{cc}\frac{v_{\chi;+}}{\sqrt{2}}&0\\ v_{\chi;0}&-\frac{v_{\chi;+}}{\sqrt{2}}\end{array}\right),~{}\xi=\left(\begin{array}[]{cc}\frac{v_{\xi;0}}{\sqrt{2}}&-v_{\xi;+}\\ -v_{\xi;+}&-\frac{v_{\xi;0}}{\sqrt{2}}\end{array}\right).$ (97)
@@
    $V(v_{\phi},v_{\chi;+},v_{\chi;0},v_{\xi;+},v_{\xi;0})$ $=$ $\frac{\lambda}{4}v_{\phi}^{4}+\rho_{1}\left(v_{\chi;+}^{2}+v_{\chi;0}^{2}\right)^{2}+\rho_{2}\left(\frac{1}{2}v_{\chi;+}^{4}+v_{\chi;0}^{4}+3v_{\chi;+}^{2}v_{\chi;0}^{2}\right)$ (98)
@@
    It should be noted that the couplings, such as $\lambda_{i},\rho_{i},\sigma_{i}$, etc., must be replaced with their values in regions of large field values evolved by renormalization group. These evolved values can be utilized to determine the vacuum structure of the GM model. We disregard the dimensional parameters, which can be eliminated through the minimization conditions for the VEVs and play a negligible role in deriving the positive definiteness constraints in regions of large field values for the effective scalar potential.
    In the large field value approximation, the VEVs of preserving custodial EWSB vacuum of the GM model can be estimated as
    $v_{\phi}\approx 0,~{}~{}v_{\chi;0}=v_{\xi;0}\approx 0,~{}~{}v_{\chi;+}=v_{\xi;+}=0~{}.$ (99)
    The constraints of vacuum stability demand that no deeper vacua, with the exception of the ordinary custodial symmetry-preserving vacuum, can emerge in regions of large field values. In other words, it is required that the potential $V(v_{\phi},v_{\chi;+},v_{\chi;0},v_{\xi;+},v_{\xi;0})$ be positive definite.

### v2

@@ 1 Introduction
$\mu_{\gamma\gamma}^{ATLAS+CMS}={0.24}_{-0.08}^{+0.09}.$ (1)
@@
Given such interesting coincidental signals, it is meaningful to explore whether the final result corresponds to a new light scalar particle from the perspective of new physics. Indeed, many papers have explained these signals within different new physics models combine:95 ; 95-1 ; 95-2 ; 95-3 ; 95-4 ; 95-5 ; 95-6 ; 95-7 ; 95-8 ; 95-9 ; 95-10 ; 95-11 ; 95-12 ; 95-13 ; 95-14 ; 95-15 ; 95-16 ; 95-17 ; 95-18 ; 95-19 ; 95-20 ; 95-21 ; 95-22 ; 95-23 ; 95-24 ; 95-25 ; 95-26 ; 95-27 ; 95-28 ; 95-29 ; 95-30 ; 95-31 ; 95-32 ; 95-33 ; 95-34 ; 95-35 ; 95-36 ; 95-37 ; 95-38 ; 95-39 ; 95-40 ; 95-41 ; 95-42 ; 95-43 ; 95-44 ; 95-45 ; 95-46 ; 95-47 ; 95-48 ; 95-49 ; 95-50 ; 95-51 ; 95-52 ; 95-53 ; 95-54 ; 95-55 ; 95-56 ; 95-57 ; 95-58 ; 95-59 ; 95-60 ; 95-61 ; 95-62 ; 95-63 ; 95-64 ; 95-GM-1 ; 95-GM-2 ; 95-GM-3 .
@@
Among these models, the GM model can provide a light custodial singlet scalar as a perfect candidate to interpret di-photon and $b\bar{b}$ excess simultaneously 95-GM-1 ; 95-GM-2 ; 95-GM-3 . There could even be a twin-peak resonance around 95 GeV 95-GM-2 , and demonstrates the nature of di-photon, $b\bar{b}$, and di-tau excesses with two scalars in the GM model (a light $CP$-even scalar and a same-mass $CP$-odd scalar).
@@
The GM model  95-GM-1 ; 95-GM-2 ; 95-GM-3 ; GM ; GM2 ; GM3 ; GM4 ; GM5 ; GM6 ; GM7 ; GM8 ; GM9 ; GM10 ; GM11 ; GM12 ; GM13 ; GM14 ; GM15 ; GM16 ; GM17 ; GM18 ; GM19 ; GM20 ; GM21 ; GM22 ; GM23 ; GM24 ; GM25 ; GM26 ; GM27 ; GM28 ; GM29 ; GM30 ; GM31 ; GM32 ; GM33 ; GM34 ; GM35 ; GM36 ; GM37 ; GM38 ; GM39 ; GM40 ; GM41 ; GM63 ; GM64 ; GM66 ; GM68 ; GM69 ; GM70 ; GM44 ; GM45 ; GM46 ; GM47 ; GM48 ; GM49 ; GM50 ; GM51 ; GM52 ; GM53 ; GM54 ; GM55 ; GM56 ; GM57 ; GM58 ; GM59 ; GM60 ; GM61 ; GM62 ; GM65 ; GM67 ; GM-pd extends the SM by incorporating a complex $SU(2)_{L}$ triplet scalar $\chi$ with unit hypercharge and a real $SU(2)_{L}$ triplet scalar $\xi$ with zero hypercharge. This intricate construction enables the nontrivial preservation of custodial symmetry at tree level after EWSB, dependent on the precise alignment of the vacuum expectation values (VEVs) between these triplet scalars, which ensures that the triplet state expectation value can reach several tens of GeV, significantly higher than in new physics models that extend the SM with a triplet. In addition to the unique additional scalar sector, the GM model predicts interesting phenomenology GM24 ; GM25 ; GM26 ; GM27 ; GM28 ; GM29 ; GM30 ; GM31 ; GM32 ; GM33 ; GM34 ; GM35 ; GM36 ; GM37 ; GM38 ; GM39 ; GM40 ; GM41 ; GM63 ; GM64 ; GM66 , such as enhanced couplings of the SM-like Higgs boson to W and Z gauge bosons, as well as the existence of nine more physical scalars, including a doubly charged scalar particle.
@@
In the GM model related studies, theoretical constraints from perturbative unitarity, the EW vacuum stability and conditions to avoid alternative minima vacuum GM57 ; GM58 ; GM59 ; GM60 ; GM61 ; GM62 ; GM65 ; GM67 etc should be satisfied other than those exclusion bounds from new particle searches GM56 ; GM59 by colliders. In our previous work GM-pd , we have rediscussed the vacuum stability bound for the GM model, concentrating specially on the constraints bounds from non-existence of a deeper vacuum at large field values. The numerical results indicate that quite a lot parameter space satisfied theoretical and experimental constraints have been excluded by the new vacuum stability bounds. It is interesting to explore whether there is still parameter space within the GM model that can accommodate the 95 GeV excess, considering all the theoretical and experimental constraints.
@@ 2 A brief review of the Georgi-Machacek model
The GM model’s Higgs sector encompasses the well-known SM-like $SU(2)_{L}$ complex doublet Higgs field, labeled as $({\phi}^{+},~{\phi}^{0})$, which carries a hypercharge of 1/21/21 / 2. Furthermore, it incorporates two additional $SU(2)_{L}$ triplet Higgs fields: one complex triplet, identified as $({\chi}^{++},~{\chi}^{+},~{\chi}^{0})$) with a unit hypercharge, and a real triplet possessing a zero hypercharge, denoted as $({\xi}^{+},~{\xi}^{0},~-{\xi}^{+*})$. This extension introduces novel features and complexities in the Higgs sector compared to the Standard Model. The general potential [^2] The last two terms in the scalar potential can be absent if introducing a $Z_{2}$ symmetry. of GM model being invariant under the global $SU(2)_{L}\times SU(2)_{R}$ symmetry and SM gauge symmetry can be given as
@@
$m_{h,H}^{2}=\frac{\mathcal{M}_{11}^{2}+\mathcal{M}_{22}^{2}\mp\sqrt{{\left(\mathcal{M}_{11}^{2}-\mathcal{M}_{22}^{2}\right)}^{2}+4{\left(\mathcal{M}_{12}^{2}\right)^{2}}}}{2}.$ (48)
@@ 3.1 Brief review of the 95 GeV excess
$\mu_{\gamma\gamma}^{ATLAS+CMS}\equiv\frac{\sigma^{exp}(gg\rightarrow\phi\rightarrow\gamma\gamma)}{\sigma^{SM}(gg\rightarrow h\rightarrow\gamma\gamma)}={0.24}_{-0.08}^{+0.09}.$ (51)
@@ 3.2 The 95 GeV excess at the GM model
$\mu(\Phi_{95\mathrm{GeV}})\equiv\frac{\sigma({\Phi_{95\mathrm{GeV}}})\times BR(\Phi_{95\mathrm{GeV}}\rightarrow\gamma\gamma/b\bar{b})}{\sigma_{\mathrm{SM}}({\Phi_{95\mathrm{GeV}}})\times BR_{\mathrm{SM}}(\Phi_{95\mathrm{GeV}}\rightarrow\gamma\gamma/b\bar{b})}.$ (53)
@@
$\kappa_{Hff}=g_{H_{\mathrm{{GM}}ff}}/g_{h_{\mathrm{SM}}ff},~~~\kappa_{HVV}=g_{H_{\mathrm{GM}}VV}/g_{h_{\mathrm{SM}}VV}.$ (55)
In the investigation of resonances around 95 GeV, the choice for $ff$ should be $t\bar{t}$, and for $VV$, it should be $ZZ$ (these choices are due to the requirements of the corresponding resonance production modes at both LHC and LEP). These selections correspond to the production modes that are relevant to both LEP and LHC experiments. For the GM model, the effective Higgs couplings to $t\bar{t}$ and $ZZ$ in the lowest order can be shown as
$\kappa_{Ht\bar{t}}=\frac{\sin{\alpha}}{\cos{\theta}},~~~\kappa_{ht\bar{t}}=\frac{\cos{\alpha}}{\cos{\theta}},$
$\kappa_{HZZ}=\cos{\theta}\sin{\alpha}+\sqrt{\frac{8}{3}}\sin{\theta}\cos{\alpha},~~~\kappa_{hZZ}=\cos{\theta}\cos{\alpha}-\sqrt{\frac{8}{3}}\sin{\theta}\sin{\alpha},$ (56)
@@ 4.1 The Usual Constraints
[…]
@@
[…]
@@
    $\kappa_{tt}=1.01\pm 0.11,~\kappa_{bb}=0.99\pm 0.16,~\kappa_{\tau\tau}=0.92\pm 0.08,$
@@
    $0.83\leq\kappa_{tt}\leq 1.10,~0.78\leq\kappa_{bb}\leq 1.15,~0.86\leq\kappa_{\tau\tau}\leq 1.00,$
@@
    $1.7\times 10^{-9}<BR(B^{0}_{s}\rightarrow\mu^{+}\mu^{-})<4.5\times 10^{-9}~,$
    $0.85\times 10^{-4}<BR(B\rightarrow\tau\nu)<2.89\times 10^{-4}~,$
    $2.99\times 10^{-4}<BR(B\rightarrow X_{s}\gamma)<3.87\times 10^{-4}~.$ (76)
@@ 4.2 Positive Definiteness Constraints
In light of these considerations, we have introduced the so-called “Positive Definiteness Constraints” within the GM model to ensure the stability of the custodial vacuum GM-pd . Here we briefly review the Positive Definiteness Constraints on the GM model.
@@
    $+\mu^{\prime}_{1}\phi^{\dagger}\xi\phi+\mu^{\prime}_{2}[\phi^{T}(i\tau_{2})\chi^{\dagger}\phi+\text{h.c.}]+\mu^{\prime}_{3}\text{tr}(\chi^{\dagger}\chi\xi)$
@@
    $m_{\phi}^{2}$ $=$ $2m_{\Phi}^{2},~m_{\chi}^{2}=2m_{\Delta}^{2},~m_{\xi}^{2}=m_{\Delta}^{2},~\mu^{\prime}_{1}=-\frac{\mu_{1}}{\sqrt{2}},~$
    $\mu^{\prime}_{2}$ $=$ $-\frac{\mu_{1}}{2},~\mu^{\prime}_{3}=6\sqrt{2}\mu_{2},\lambda=4\lambda_{1},~\rho_{1}=4\lambda_{2}+6\lambda_{3},~$
@@
    A comprehensive analysis of the effects of custodial symmetry breaking in the GM model at higher energies can be found in Ref. GM44 , assuming that no custodial symmetry-violating terms are present at the EW scale. The study demonstrates that, within certain parameter regions, the influence of custodial symmetry breaking can be adequately controlled up to the grand unified theory (GUT) scale. Alternatively, if custodial symmetry is preserved at high-energy scales within a UV completion of the GM model, the measured value of the electroweak $\rho$ parameter, along with the condition of perturbative unitarity, places stringent limits on the scale of this ultraviolet completion across almost the entire parameter space GM45 .
    Despite their different approaches, the two methods used to explore the custodial symmetry breaking effects induced by RGE between various energy scales yield equivalent consequences in the frame of the GM model. And one-loop Renormalization Group Equations (RGEs) for scalar quartic couplings $\lambda,~\rho_{1,2,3,4,5},~\sigma_{1,2,3,4}$, the gauge couplings $g_{i}(i=1,2,3)$, and the Yukawa couplings $Y_{t,b,\tau}$, etc. have been provided in GM44 ; GM45 , with two-loop RGEs also available in GM65 , which can be used to conduct the RGE analysis. In practical numerical analyses, the Mathematica package SARAH SARAH1 ; SARAH2 ; SARAH3 ; SARAH4 ; SARAH5 not only facilitates the generation of source codes for the spectrum-generator package SPheno SPheno1 ; SPheno2 but also can be utilized for RGE numerical studies.
@@
    The invariance of the $SU(2)_{L}$ gauge symmetry allows us to rotate away three redundant VEVs of the four components of weak isospin components, leaving only $v_{\phi}$. Furthermore, the invariance of the $SU(2)_{V}$ custodial symmetry can be used to render three of the VEVs within the triplets redundant, so we can choose to eliminate the VEVs for $\chi^{++}$. Besides, considering any attempt of spontaneous $CP$ breaking would lead to the breaking of the custodial symmetry at EW scale GM60 , we can assume $CP$ conservation ($v_{\chi;+}^{I}=v_{\chi;0}^{I}=0$) in this work. We are left with the following form of the VEVs after denoting $v_{\chi;+}^{R}\equiv v_{\chi;+}$ and $v_{\chi;0}^{R}=v_{\chi;0}$ for simplicity
    $\langle\phi\rangle=\left(\begin{array}[]{c}0\\ \frac{v_{\phi}}{\sqrt{2}}\end{array}\right),~\langle\chi\rangle=\left(\begin{array}[]{cc}\frac{v_{\chi;+}}{\sqrt{2}}&0\\ v_{\chi;0}&-\frac{v_{\chi;+}}{\sqrt{2}}\end{array}\right),~\langle\xi\rangle=\left(\begin{array}[]{cc}\frac{v_{\xi;0}}{\sqrt{2}}&-v_{\xi;+}\\ -v_{\xi;+}&-\frac{v_{\xi;0}}{\sqrt{2}}\end{array}\right).$ (93)
@@
    $V(v_{\phi},v_{\chi},v_{\xi})$ $=$ $\frac{\lambda}{4}v_{\phi}^{4}+\rho_{1}\left(v_{\chi;+}^{2}+v_{\chi;0}^{2}\right)^{2}+\rho_{2}\left(\frac{1}{2}v_{\chi;+}^{4}+v_{\chi;0}^{4}+3v_{\chi;+}^{2}v_{\chi;0}^{2}\right)$ (94)
@@
    It should be noted that the couplings, such as $\lambda,\rho_{i},\sigma_{i}$, etc., must be replaced with their values in regions of large field values evolved by RGEs. These evolved values can be utilized to determine the vacuum structure of the GM model. We disregard the dimensional parameters, which can be eliminated through the minimization conditions for the VEVs and play a negligible role in deriving the positive definiteness constraints in regions of large field values for the effective scalar potential.
    It is evident that the VEVs of the custodial-preserving EWSB vacuum are significantly smaller than those of potential new vacua which may emerge in regions of large field values. As a consequence, the magnitude of the potential energy density for the custodial-preserving EWSB vacuum is considerably lower than that of the deeper vacua which arise at large field values, with the exception of certain highly exceptional cases. To ensure vacuum stability, it is imperative that no lower-lying vacua, other than the conventional custodial-preserving vacuum, develop in regions of large field values. In other words, the scalar potential $V(v_{\phi},v_{\chi},v_{\xi})$ must remain positive definite across all field-space asymptotes, analogous to the setting of ordinary vacuum stability bounds in the Standard Model.
@@
    It should be noted that, while we have utilized the one-loop RG-improved tree-level scalar potential to numerically analyze the vacuum structure and the constraints of positive definiteness in the GM model, we have not considered the possibility that high-energy new physics could potentially destabilize the vacuum, leading to rapid vacuum decay with a tunneling time shorter than the age of the universe rvdthnp-1 ; rvdthnp-2 ; rvdthnp-3 ; rvdthnp-4 ; rvdthnp-5 . Moreover, the inclusion of gravitational effects at high energies could also significantly impact the decay rate of the false vacuum rvdthnp-5 ; rvdthnp-g-1 ; rvdthnp-g-2 . These aspects will be thoroughly investigated in our future works.